# 3cm3cm2cm2cm

### v1

@@ Abstract
Evaluating the $H(x,s|pp)$ scaling function of elastic proton-proton ($pp$) collisions from recent TOTEM data at $\sqrt{s}=8$ TeV and comparing it with the same function of elastic proton-antiproton ($p\bar{p}$) data of the D0 collaboration at $\sqrt{s}=1.96$ TeV, we find, from this comparison alone, an at least 3.79 $\sigma$ signal of Odderon exchange. If we combine this model independently obtained result with that of a similar analysis but using TOTEM elastic $pp$ scattering data at $\sqrt{s}=7$ TeV, which resulted in an at least 6.26 $\sigma$ signal, the combined significance of Odderon exchange increases to at least 7.08 $\sigma$, model independently. Further combinations of various datasets in the TeV energy range are detailed in the manuscript.
@@ 1 Introduction
A direct way to probe odderon exchange in elastic scattering is by comparing the differential cross-section of particle-particle and particle-antiparticle scattering at the same and sufficiently high energy [3, 4]. A search performed at the ISR energy of $\sqrt{s}=53$ GeV in 1985 [5] resulted in an indication of the Odderon at the 3.35 $\sigma$ significance level. That analysis, however, did not utilize all the available data in the overlapping acceptance of the $pp$ and $p\bar{p}$ measurements. Furthermore, at $\sqrt{s}=53$ GeV, Reggeon exchanges may play a significant role, rendering the Odderon search at the ISR energies rather inconclusive.
@@
As far as we know, the first anonymously peer-reviewed publication of a statistically significant, at least 6.26 $\sigma$ signal of odderon exchange was published by the authors of the present manuscript, in the proceedings of the ISMD 2019 (Santa Fe, NM, USA), published in its final version in June 2020 [6]. This refereed conference contribution was backed up and detailed in our publication, published in its final form in February 2021, with the same statistical significance of at least 6.26 $\sigma$ signal for odderon exchange [7]. Our Hungarian-Swedish team introduced a new scaling function that turned out to be energy independent in the LHC energies of $\sqrt{s}=2.76-7.0$ TeV in elastic proton-proton ($pp$) collisions, based on a model-independent, direct data to data comparison [7].
@@
These results as well as the model-independent investigation of the domain of validity of the $H(x,s|pp)$ scaling have been seconded by a theoretical paper of T. Csörgő, and I. Szanyi [8], increasing the statistical significance of the observation of odderon exchange to at least 7.08 $\sigma$. At the same time, this model-dependent investigation found that the domain of validity in $s$ of the $H(x,s|pp)=H(x,s_{0}|pp)$ scaling (where $s_{0}$ stands for a fixed energy scale in the TeV region, e.g. $\sqrt{s_{0}}=7$ TeV) extends also to the top Tevatron energies of $\sqrt{s}=1.96$ TeV [8]. This theoretical work utilized a validated model, proposed in its first form by A. Bialas and A. Bzdak [9], however, the original model lacked a real part hence the possibility of odderon exchange. However, the so-called Real-extended Bialas-Bzdak (ReBB) model of Ref. [10] fixed these shortcomings and has been utilized in Ref. [8] to extrapolate not only the elastic proton–proton scattering data from the LHC energies of $\sqrt{s}=2.76$ and 7 TeV to the D0 energy of $\sqrt{s}=1.96$ TeV but also to extrapolate the elastic proton–antiproton scattering data from $\sqrt{s}=0.546$ and 1.96 TeV to the LHC energies of 2.76 TeV and 7 TeV. Evaluating the proton–proton data with a model increased the uncertainty and decreased the odderon signal from proton–proton scattering data alone, but this decrease was well over-compensated with the ability of the model to evaluate theoretically the proton–antiproton scattering at all the LHC energies. Overall, this procedure resulted in a model-dependent increase of the statistical significance from odderon exchange from 6.26 to 7.08 $\sigma$ [8] as published in final form in July 2021, but limited the comparison to the diffractive minimum and maximum region in the four-momentum transfer range, to the domain of verified validity of the ReBB model. More recently, these results were extended to the new TOTEM data on elastic $pp$ scattering at $\sqrt{s}=8$ TeV, published in March 2022 [11] in Ref. [8]. When TOTEM data on elastic $pp$ collisions at $\sqrt{s}=8$, 7, and 2.76 TeV are analyzed simultaneously with D0 data at $\sqrt{s}=$ 1.96 TeV in the framework of the ReBB model, a combined statistical significance greater than 32.4 $\sigma$ can be achieved, rendering the statistical significance of odderon observation, in any practical terms, to a certainty [8].
@@
These studies utilized a series of important papers published by the TOTEM Collaboration investigating the properties of elastic $pp$ scattering in the LHC energy range between $\sqrt{s}=2.76$ and $13$ TeV [12, 13, 14, 15]. Most recently, the latest measurement performed by TOTEM at $\sqrt{s}=8$ TeV [11] extended the earlier analysis up to $|t|$-values of 1.9 GeV2 [16]. An increase of the total cross section, $\sigma_{\rm tot}(s)$, associated with a decrease of the real-to-imaginary ratio, $\rho(s)$, with energy, first identified at $\sqrt{s}=13$ TeV [12, 13] indicated a possible Odderon effect triggering an intense discussion in the literature (see e.g. Refs. [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]). The persistent diffractive minimum-maximum structure in the $t$-dependent profile of $d\sigma/dt$ in elastic $pp$ collisions observed by TOTEM at $\sqrt{s}$ = 2.76, 7, 8 and 13 TeV, and the lack of such structure in elastic $p\bar{p}$ collisions measured by D0 [35] at $\sqrt{s}=1.96$ TeV, indicate a qualitatively clear Odderon effect. The possibility of utilizing experimental data at the TeV energy scale for a search for a statistically significant odderon exchange has been proposed by Jenkovszky et al. in Refs. [4, 19]. In 2020, the TOTEM collaboration made a qualitative conclusion about odderon exchange in Ref. [15] as follows: “Under the condition that the effects due to the energy difference between TOTEM and D0 can be neglected, the result provides evidence for a colourless 3-gluon bound state exchange in the $t$-channel of the $pp$ elastic scattering”. However, no statistical significance for this observation has been evaluated in Ref. [15].
@@
More recently, in August 2021, a properly quantified statistical significance of the Odderon signal has been published by the TOTEM and D0 Collaborations [36] employing different methods and techniques, obtaining an at least 5.2 $\sigma$ combined statistical significance for an almost model independent observation of odderon exchange, a first statistically significant result obtained by two experimental collaborations. This result was based on the extrapolation of TOTEM experimental data of the differential cross-section of elastic $pp$ scattering from $\sqrt{s}=$ 13, 8, 7 and 2.76 TeV down to 1.96 TeV using an almost model independent technique and comparing the results with D0 data in a limited four-momentum transfer range, resulting in an at least 3.4 $\sigma$ signal for Odderon exchange. TOTEM also has measured the pair of the total cross-section and the real-to-imaginary ratio $(\sigma_{tot},\rho)$ and compared it with a set of models without odderon exchange. When a partial combination of the TOTEM $(\sigma_{tot},\rho)$ measurements is done at $\sqrt{s}=13$ TeV, the obtained partial significances range between 3.4 and 4.6 $\sigma$ for the considered models. The full combination of the signal of odderon exchange from TOTEM $(\sigma_{tot},\rho)$ measurements at $\sqrt{s}=13$ TeV with the signal of the comparision of extrapolated TOTEM $pp$ data to $\sqrt{s}=1.96$ TeV with a subset of 8 out of 17 D0 datapoints [37] on elastic $p\bar{p}$ scattering leads to total significances ranging from 5.2 to 5.7 $\sigma$ for $t$-channel odderon exchange for each of the considered models [36].
@@
The validity of this D0-TOTEM proof of Odderon exchange has been questioned in several published papers by now. Most importantly, the ATLAS collaboration [38] published a significantly different total cross-section hence a significantly different pair of $(\sigma_{tot},\rho)$ at $\sqrt{s}=13$ TeV, questioning the significance of the signal of odderon exchange from these low −$-t$ observations. Such a significant incompatibility between the ATLAS and TOTEM measurements of total cross-section and the ratio of real to imaginary part of the scattering amplitude, that is between the pairs of $(\sigma_{tot},\rho)$ at $\sqrt{s}=13$ TeV as published by the ATLAS and by the TOTEM experiments has recently been confirmed by an independent analysis of Petrov and Tkachenko in Ref. [39]. Furthermore, Donnachie and Landshoff [40, 41] stressed the point that phase of an elastic scattering amplitude is related to its energy variation, and as a consequence, they have questioned the validity of the D0-TOTEM signal of odderon exchange at $t=0$. Petrov and Tkachenko obtained results similar to that of Donnachie and Landshoff in Ref. [42], suggesting that the systematic error on the determination of the $\rho$ parameter at $\sqrt{s}=13$ TeV might have been slightly but significantly underestimated by TOTEM in Ref. [43]. Let us mention that Refs. [6, 7] scale out the $t=0$ observables from their analysis of the $H(x,s|pp)$ scaling functions, while the low-$t$ domain has been explicitely excluded from finding a statistically significant signal of odderon-exchange in Refs. [44, 8]. Hence these odderon discovery papers are not affected by the above mentioned criticisms of the D0-TOTEM publication of a statistically significant, at least 5.2 $\sigma$ experimental observation of odderon exchange.
@@
A response to these published criticisms was given by the talk of K. Österberg, the physics coordinator of the TOTEM experiment at the ISMD 2023 conference in Gyöngyös, Hungary. We have good reasons to expect that a detailed D0-TOTEM response to the above criticisms will be submitted for a publication as soon as reasonably possible. Furthermore, we also second the suggestion of Petrov and Tkachenko in Ref. [39], calling for a joint ATLAS-TOTEM analysis to sort out the differences between their low-$t$ measurements at $\sqrt{s}=13$ TeV, proposing also a comparison of ATLAS and TOTEM data at low −$-t$ at $\sqrt{s}=7$ TeV [46, 47, 48] and the same comparison also at $8$ TeV [49, 50].
@@
Our approach is model independent in the sense that it does not rely on any fitting function or any theoretical input, it uses only linear and log-linear interpolation techniques between neighbouring datapoints, to allow for data to data comparison at the same values of the horizontal axes (using the scaling variable $x=-tB$). As we compare pairwise the $H(x,s|pp)$ scaling functions constructed at different energies based only upon the available data and look for statistically significant differences within any pair of TeV-scale $pp$ and $p\bar{p}$ data sets depending on the collision energy, we need to utilize rebinning that also includes model independent linear interpolation methods, as detailed in Ref. [7], where the basic concepts and methodology have been explained in detail.
@@
A similar function $H(x,s|p\bar{p})$ can be introduced for elastic proton-antiproton ($p\bar{p})$ collisions. In Ref. [7] we have shown, that $H(x,s|pp)=H(x|pp)$ is independent of the energy in elastic $pp$ collisions, utilizing direct data to data comparison of elastic $pp$ data of the TOTEM Collaboration at $\sqrt{s}=2.76$ and $7$ TeV, as described in Refs. [53, 54, 46]. In Ref. [7] we have also shown, in a model-dependent way, that this scaling extends down to $\sqrt{s}=1.96$ TeV, and we have also found that the scaling violations are significant already at $\sqrt{s}=13$ TeV. In the present study we investigate if the energy independence of this scaling is, within experimental errors, valid for the recently published TOTEM data at $\sqrt{s}=8$ TeV, or not, within the experimental errors as determined by the TOTEM collaboration in Ref.  [11]. This question is particularly interesting as the elastic to total cross-section ratio increases with increasing energies, and at the LHC energies between 8 and 13 TeV it crosses significantly the important limit of $(1+\rho^{2})\sigma_{\rm el}/\sigma_{\rm tot}=1/4$, see Fig. 4 of Ref. [55]. This ratio reaches such a critical value in the region of $\sqrt{s}=2.76$ – $7.0$ TeV and it clearly exceeds it at $\sqrt{s}=13.0$ TeV, where we find statistically significant violations of the $H(x,s|pp)=H(x,s_{0}|pp)$ scaling law. The search for the onset of the violation of $H(x,s|pp)=H(x,s_{0}|pp)=H(x,pp)$ scaling motivates our current investigation.
@@
In the left panels of Fig. 1 and Fig. 2, we demonstrate the energy independence of the $H(x|pp)$-scaling within statistical errors, where the TOTEM datasets at $\sqrt{s}=2.76$ and $8$ TeV as well as at $\sqrt{s}=7$ and $8$ TeV are compared pairwise, respectively. The agreement between these datasets corresponds to a confidence level (CL) of at least 98% . Thus such a $H(x)$-scaling removes the trivial energy dependent terms, due to the known $s$-dependence of the elastic slope $B(s)$, the elastic and total cross-sections $\sigma_{\rm el}(s)$ and $\sigma_{\rm tot}(s)$, and the real-to-imaginary ratio $\rho(s)$ [55].
However, one should note that in a local low $x$ (low $|t|$) interval one can see larger deviations than the reported data errors suggest: In case of the comparison of the $H(x,s|pp)$ scaling functions at $\sqrt{s}=7$ TeV and $8$ TeV, we observe a partial disagreement beyond the published experimental errors in a small $x$ range of about $|t|=[0.175-0.375]$ GeV2, which points toward a possible adjustment problem of the two parts of the data sets measured with different optics ($\beta^{*}=3.5$ m and $\beta^{*}=90$ m and published separately in Tables 5 and 4 of Ref. [46], respectively) and suggests the need of a further, more detailed investigation of the published systematic errors in the low $|t|=[0.175-0.375]$ GeV2 range of the dataset measured by TOTEM at $\sqrt{s}=7$ TeV with the $\beta^{*}=90$ m LHC optics [46].
@@
The $8\to 1.96$ TeV and $1.96\to 8$ TeV projections correspond to keeping the measured $x$ values at $\sqrt{s}=$ 1.96 and 8 TeV, respectively, and determining by interpolation the $H(x,s)$ scaling functions at these $x$ values, but at the other energy, $\sqrt{s}=$ 8 and 1.96 TeV, respectively. This procedure is described in full details in Ref. [7]. The left panel of Fig. 3 compares the $H(x)$ scaling function of elastic $pp$ collisions at $\sqrt{s}=8$ TeV to that of $p\bar{p}$ collisions at $\sqrt{s}=1.96$ TeV. In this case, adopting the method of Ref. [56], the confidence level of the agreement of the $H(x)$ scaling functions is found to be lower then $0.015$%, with a minimum of $\chi^{2}/{\rm NDF}=46.4/17$. Hence, the difference between these scaling functions is statistically significant and represents our result for the Odderon observation from the comparison of the $\sqrt{s}=8$ TeV $pp$ and $\sqrt{s}=1.96$ TeV $p\bar{p}$ datasets to be an at least 3.79 $\sigma$ effect, in the $5\leq x\lessapprox 20$ acceptance domain. This seems to be a conservative and robust result as we find that this value can only increase for variations both in the procedure itself and in the $\chi^{2}$ definition.

### v3

@@ Abstract
Evaluating the $H(x,s|pp)$ scaling function of elastic proton-proton ($pp$) collisions from recent TOTEM data at $\sqrt{s}=8$ TeV and comparing it with the same function of elastic proton-antiproton ($p\bar{p}$) data of the D0 collaboration at $\sqrt{s}=1.96$ TeV, we find, from this comparison alone, an at least 3.79 $\sigma$ signal of odderon exchange. If we combine this model-independently obtained result with that of a similar analysis but using TOTEM elastic $pp$ scattering data at $\sqrt{s}=7$ TeV, which resulted in an at least 6.26 $\sigma$ signal, the combined significance of odderon exchange increases to at least 7.08 $\sigma$. Further combinations of various datasets in the TeV energy range are detailed in the manuscript.
@@ 1 Introduction
A direct way to probe odderon exchange in elastic scattering is by comparing the differential cross-section of particle-particle and particle-antiparticle scattering at the same and sufficiently high energy [3, 4]. The TeV energy scale is high enough to render the amplitudes for the massive (Reggeon) exchange to negligibly small values. The differential cross-sections for elastic proton-proton as well as proton-antiproton scatterings are proportional to the modulus squares of the corresponding amplitudes. If the modulus squared amplitudes are different, then the two amplitudes cannot be equal. As the amplitude for odderon exchange is proportional to the difference between the amplitudes of the proton-antiproton and the proton-proton elastic scattering, it follows, that the amplitude for odderon exchange cannot vanish if the differential cross-sections of elastic proton-proton and proton-antiproton scatterings are significantly different at the TeV energy scale, as detailed in Ref. [5]. A search performed at the ISR energy of $\sqrt{s}=53$ GeV in 1985 [6] resulted in an indication of the odderon at the 3.35 $\sigma$ significance level. That analysis, however, did not utilize all the available data in the overlapping acceptance of the $pp$ and $p\bar{p}$ measurements. Furthermore, at $\sqrt{s}=53$ GeV, Reggeon exchanges may play a significant role, rendering the odderon search at the ISR energies rather inconclusive.
@@
As far as we know, the first anonymously peer-reviewed publication of a statistically significant, at least 6.26 $\sigma$ signal of odderon exchange was published in June 2020 by the authors of the present manuscript, in the proceedings of the ISMD 2019 (Santa Fe, NM, USA) [7]. This refereed conference contribution was backed up in our February 2021 paper, with the same statistical significance of at least 6.26 $\sigma$ signal for odderon exchange [5]. Our Hungarian-Swedish team introduced a new scaling function that turned out to be energy independent in the LHC energies of $\sqrt{s}=2.76-7.0$ TeV in elastic proton-proton ($pp$) collisions, based on a model-independent, direct data to data comparison [5]. Our idea was based on the existence of a diffraction cone at small values of the squared four-momentum transfer, the Mandelstam variable $t$, characterized by an exponential fall. By plotting the $pp$ and $p\bar{p}$ data as a function of a dimensionless variables $x=-B(s)t$, where $B$ is the characteristic scale of this exponential fall, and by dividing the differential cross-section by $B\,\sigma_{\rm el}$, where $\sigma_{\rm el}$ is the integral of the differential cross-section, (dominated by the contribution of the diffractive cone), a dimensionless $H(x,s)$ scaling function is obtained. In the region of the diffractive cone, this scaling function is approximately a structureless exponential, $H(x)\approx\exp(-x)$. As the diffractive cone is located at small values of $t$, this $H(x)$ scaling removes the trivial $s$-dependences of the differential cross-section, where the Mandelstam variable $s$ is the squared center of mass energy. Thus such a $H(x)$-scaling removes the trivial energy dependent terms, due to the known $s$-dependence of the elastic slope $B(s)$, the elastic and total cross-sections $\sigma_{\rm el}(s)$ and $\sigma_{\rm tot}(s)$, and the real-to-imaginary ratio $\rho(s)$ [8]. It follows, that the amplitude for odderon exchange cannot vanish if the dimensionless $H(x)$ scaling functions of elastic proton-proton and proton-antiproton scatterings are significantly different at the TeV energy scale  [5].
@@
These results as well as the model-dependent investigation of the domain of validity of the $H(x,s|pp)$ scaling have been seconded by a theoretical paper of T. Csörgő, and I. Szanyi [9], increasing the statistical significance of the observation of odderon exchange to at least 7.08 $\sigma$. At the same time, this model-dependent investigation found that the domain of validity in $s$ of the $H(x,s|pp)=H(x,s_{0}|pp)$ scaling (where $s_{0}$ stands for a fixed energy scale in the TeV region, e.g. $\sqrt{s_{0}}=7$ TeV) extends also to the top Tevatron energies of $\sqrt{s}=1.96$ TeV [9]. This theoretical work utilized a validated model, proposed in its first form by A. Bialas and A. Bzdak [10], however, the original model lacked a real part hence the possibility of odderon exchange. However, the so-called Real-extended Bialas-Bzdak (ReBB) model of Ref. [11] fixed these shortcomings and has been utilized in Ref. [9] to extrapolate not only the elastic proton–proton scattering data from the LHC energies of $\sqrt{s}=2.76$ and 7 TeV to the D0 energy of $\sqrt{s}=1.96$ TeV but also to extrapolate the elastic proton–antiproton scattering data from $\sqrt{s}=0.546$ and 1.96 TeV of the UA4  [12, 13] and the D0 Collaborations [14] to the LHC energies of 2.76 TeV and 7 TeV. Evaluating the proton–proton data with a model increased the uncertainty and decreased the odderon signal from proton–proton scattering data alone, but this decrease was well over-compensated with the ability of the model to evaluate theoretically the proton–antiproton scattering at all the LHC energies. Overall, this procedure resulted in a model-dependent increase of the statistical significance from odderon exchange from 6.26 to 7.08 $\sigma$ [9] as published in final form in July 2021, but limited the comparison to the diffractive minimum and maximum region in the four-momentum transfer range, to the verified domain of validity of the ReBB model. More recently, these results were extended to the new TOTEM data on elastic $pp$ scattering at $\sqrt{s}=8$ TeV, published in March 2022 [15] in Ref. [9]. When TOTEM data on elastic $pp$ collisions at $\sqrt{s}=8$, 7, and 2.76 TeV are analyzed simultaneously with D0 data at $\sqrt{s}=$ 1.96 TeV in the framework of the ReBB model, a combined statistical significance greater than 32.4 $\sigma$ can be achieved, rendering the statistical significance of odderon observation, in any practical terms, to a certainty [9].
@@
A series of important papers has been published by the TOTEM Collaboration investigating the properties of elastic $pp$ scattering in the LHC energy range between $\sqrt{s}=2.76$ and $13$ TeV in Refs. [16, 17, 18, 19]. Most recently, the latest measurement performed by TOTEM at $\sqrt{s}=8$ TeV [15] extended the earlier analysis up to $|t|$-values of 1.9 GeV2 [20]. An increase of the total cross section, $\sigma_{\rm tot}(s)$, associated with a decrease of the real-to-imaginary ratio, $\rho(s)$, with energy, first identified at $\sqrt{s}=13$ TeV [16, 17] indicated a possible odderon effect triggering an intense discussion in the literature (see e.g. Refs. [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]). The persistent diffractive minimum-maximum structure in the $t$-dependent profile of $d\sigma/dt$ in elastic $pp$ collisions observed by TOTEM at $\sqrt{s}$ = 2.76, 7, 8 and 13 TeV, and the lack of such structure in elastic $p\bar{p}$ collisions measured by D0 [14] at $\sqrt{s}=1.96$ TeV indicate a qualitatively clear odderon effect. The possibility of utilizing experimental data at the TeV energy scale for a search for a statistically significant odderon exchange has been proposed by Jenkovszky et al. in Refs. [4, 23]. In 2020, the TOTEM collaboration made a qualitative conclusion about odderon exchange in Ref. [19] as follows: “Under the condition that the effects due to the energy difference between TOTEM and D0 can be neglected, the result provides evidence for a colourless 3-gluon bound state exchange in the $t$-channel of the $pp$ elastic scattering”. However, no statistical significance for this observation has been evaluated in Ref. [19].
@@
More recently, in August 2021, a properly quantified statistical significance of the odderon signal has been published by the TOTEM and D0 Collaborations [39] employing different methods and techniques, obtaining an at least 5.2 $\sigma$ combined statistical significance for an almost model-independent observation of odderon exchange, a first statistically significant result obtained by two experimental collaborations. This result was based on the extrapolation of TOTEM experimental data of the differential cross-section of elastic $pp$ scattering from $\sqrt{s}=$ 13, 8, 7 and 2.76 TeV down to 1.96 TeV using an almost model-independent technique and comparing the results with D0 data in a limited four-momentum transfer range, resulting in an at least 3.4 $\sigma$ signal for odderon exchange. TOTEM also has measured the pair of the total cross-section and the real-to-imaginary ratio $(\sigma_{tot},\rho)$ and compared it with a set of models without odderon exchange. When a partial combination of the TOTEM $(\sigma_{tot},\rho)$ measurements is done at $\sqrt{s}=13$ TeV, the obtained partial significances range between 3.4 and 4.6 $\sigma$ for the considered models. The full combination of the signal of odderon exchange from TOTEM $(\sigma_{tot},\rho)$ measurements at $\sqrt{s}=13$ TeV with the signal of the comparision of extrapolated TOTEM $pp$ data to $\sqrt{s}=1.96$ TeV with a subset of 8 out of 17 D0 datapoints [14] on elastic $p\bar{p}$ scattering leads to total significances ranging from 5.2 to 5.7 $\sigma$ for $t$-channel odderon exchange for each of the considered models [39].
@@
The validity of this D0-TOTEM proof of odderon exchange has been questioned in several published papers by now. Most importantly, the ATLAS collaboration [40] published a significantly different total cross-section hence a significantly different pair of $(\sigma_{tot},\rho)$ at $\sqrt{s}=13$ TeV, questioning the significance of the signal of odderon exchange from these low −$-t$ observations. Such a significant incompatibility between the ATLAS and TOTEM measurements of total cross-section and the ratio of real to imaginary part of the scattering amplitude, that is between the pairs of $(\sigma_{tot},\rho)$ at $\sqrt{s}=13$ TeV as published by the ATLAS and by the TOTEM experiments has recently been confirmed by recent results of Petrov and Tkachenko in Refs. [41, 42, 43]. Furthermore, Donnachie and Landshoff [44, 45] stressed the point that phase of an elastic scattering amplitude is related to its energy variation, and as a consequence, they have questioned the validity of the D0-TOTEM signal of odderon exchange at $t=0$. Petrov and Tkachenko obtained results similar to that of Donnachie and Landshoff in Ref. [41], suggesting that the systematic error on the determination of the $\rho$ parameter at $\sqrt{s}=13$ TeV might have been slightly but significantly underestimated by TOTEM in Ref. [17]. Let us mention that Refs. [7, 5] scale out the $t=0$ observables from their analysis of the $H(x,s|pp)$ scaling functions, while the low-$t$ domain has been explicitely excluded from finding a statistically significant signal of odderon-exchange in Refs. [46, 9]. Hence these odderon discovery papers are not affected by the above mentioned criticisms of the D0-TOTEM publication of a statistically significant, at least 5.2 $\sigma$ experimental observation of odderon exchange.
Let us note, that the ATLAS data that were measured in the diffraction cone at low −$-t$ at $\sqrt{s}=$ 7, 8 and 13 TeV in Refs. [47, 48, 40], respectively, collapse to the same $H(x)$ scaling function, that is close to a structureless $H(x)\approx\exp(-x)$. Hence, the method discussed in this manuscript cannot be utilized for a search of odderon effects in these ATLAS datasets.
@@
A response to these published criticisms was given by the talk of K. Österberg, the physics coordinator of the TOTEM experiment at the ISMD 2023 conference in Gyöngyös, Hungary extending an earlier D0-TOTEM response given also by him at the 2021 Low-x Workshop at La Biodola, Italy  [50, 51]. We have good reasons to expect that a detailed D0-TOTEM response to the above criticisms will be submitted for a publication as soon as reasonably possible. Furthermore, we also second the suggestion of Petrov and Tkachenko in Ref. [43], calling for a joint ATLAS-TOTEM analysis to sort out the differences between their low-$t$ measurements at $\sqrt{s}=13$ TeV, proposing also a comparison of ATLAS and TOTEM data at low −$-t$ at $\sqrt{s}=7$ TeV [52, 53, 47] and the same comparison also at $8$ TeV [54, 48].
@@
Our approach is model-independent in the sense that it does not rely on any fitting function or any theoretical input, it uses only linear and log-linear interpolation techniques between neighbouring datapoints, to allow for data to data comparison at the same values of the horizontal axes (using the scaling variable $x=-tB$). As we compare pairwise the $H(x,s|pp)$ scaling functions constructed at different energies based only upon the available data and look for statistically significant differences within any pair of TeV-scale $pp$ and $p\bar{p}$ data sets depending on the collision energy, we need to utilize rebinning that also includes model-independent linear interpolation methods, as detailed in Ref. [5], where the basic concepts and methodology have been explained in detail.
@@
A similar function $H(x,s|p\bar{p})$ can be introduced for elastic proton-antiproton ($p\bar{p})$ collisions. In Ref. [5] we have shown, that $H(x,s|pp)=H(x|pp)$ is independent of the energy in elastic $pp$ collisions, utilizing direct data to data comparison of elastic $pp$ data of the TOTEM Collaboration at $\sqrt{s}=2.76$ and $7$ TeV, as described in Refs. [19, 56, 52]. In Ref. [5] we have also shown, in a model-dependent way, that this scaling extends down to $\sqrt{s}=1.96$ TeV, and we have also found that the scaling violations are significant already at $\sqrt{s}=13$ TeV. In the present study we investigate if the energy independence of this scaling is, within experimental errors, valid for the recently published TOTEM data at $\sqrt{s}=8$ TeV, or not, within the experimental errors as determined by the TOTEM collaboration in Ref.  [15]. This question is particularly interesting as the elastic to total cross-section ratio increases with increasing energies, and at the LHC energies between 7 and 13 TeV it crosses significantly the important limit of $(1+\rho^{2})\sigma_{\rm el}/\sigma_{\rm tot}=1/4$, see Fig. 4 of Ref. [8]. This ratio reaches such a critical value in the region of $\sqrt{s}=2.76$ – $7.0$ TeV and it clearly exceeds it at $\sqrt{s}=13.0$ TeV, where we find statistically significant violations of the $H(x,s|pp)=H(x,s_{0}|pp)$ scaling law. The search for the onset of the violation of $H(x,s|pp)=H(x,s_{0}|pp)=H(x,pp)$ scaling motivates our current investigation.
@@
In the left panels of Fig. 1 and Fig. 2, we demonstrate the energy independence of the $H(x|pp)$-scaling within statistical errors, where the TOTEM datasets at $\sqrt{s}=2.76$ and $8$ TeV as well as at $\sqrt{s}=7$ and $8$ TeV are compared pairwise, respectively. The agreement between these datasets corresponds to a confidence level (CL) of at least 98% .
However, one should note that in a local low $x$ (low $|t|$) interval one can see larger deviations than the reported data errors suggest: In case of the comparison of the $H(x,s|pp)$ scaling functions at $\sqrt{s}=7$ TeV and $8$ TeV, we observe a partial disagreement beyond the published experimental errors in a small $x$ range of about $x=[3.48-7.45]$ on Fig. 2 corresponding to $|t|=[0.175-0.375]$ GeV2, which points toward a possible adjustment problem of the two parts of the data sets measured with different optics ($\beta^{*}=3.5$ m and $\beta^{*}=90$ m and published separately in Tables 5 and 4 of Ref. [52], respectively) and suggests the need of a further, more detailed investigation of the published systematic errors in the low $|t|=[0.175-0.375]$ GeV2 range of the dataset measured by TOTEM at $\sqrt{s}=7$ TeV with the $\beta^{*}=90$ m LHC optics [52].
@@
The $8\to 1.96$ TeV and $1.96\to 8$ TeV projections correspond to keeping the measured $x$ values at $\sqrt{s}=$ 1.96 and 8 TeV, respectively, and determining by interpolation the $H(x,s)$ scaling functions at these $x$ values, but at the other energy, $\sqrt{s}=$ 8 and 1.96 TeV, respectively. This procedure is described in full details in Ref. [5].